### v2

Abstract
[…]
LoRA enables parameter-efficient specialization of large language models through modular adapters, resulting in rapidly growing public adapter pools spanning diverse tasks. Effectively using these adapters requires routing: selecting and composing the appropriate adapters for a query. We introduce LoRAuter, a novel routing framework that selects and composes low-rank adaptation (LoRA) adapters using task representations rather than adapter characteristics. Unlike existing approaches that map queries directly to adapters, LoRAuter routes queries via task embeddings derived from small validation sets and does not require adapter training data. By operating at the task level, LoRAuter achieves efficient routing that scales with the number of tasks rather than the number of adapters. Experiments across multiple tasks show that LoRAuter consistently outperforms baseline routing approaches, matching Oracle performance (101.2101.2%) when task-aligned adapters exist and achieving state-of-the-art results on unseen tasks (+$5.2$ points). We further demonstrate the robustness of LoRAuter to very large, noisy adapter pools by scaling it to over $1500$ adapters.
1 Introduction
[…]
Effectively using these adapter pools during inference introduces a critical bottleneck: adapter routing. Given a user query, how do we efficiently identify and select an appropriate subset of adapters? Existing approaches for LoRA adapter routing formulate this as a granular retrieval problem, attempting to map each individual query directly to a small number of adapters, typically between one and four. However, this strategy faces several limitations. First, existing methods either rely on a static pool of adapters Huang et al. (2024); Wu et al. (2024) or require access to the original adapter training data to construct a searchable index Zhao et al. (2024); Chronopoulou et al. (2023), conditions that rarely hold for dynamic, public adapter pools where training data is proprietary or withheld for privacy reasons. Second, methods that overcome this issue by directly routing based on adapter parameter values Ostapenko et al. (2024) incur routing overheads that scale linearly with the number of adapters and layers, making them expensive for large libraries. Finally, formulating routing as a direct matching problem between queries and adapters fails to exploit the inherent structure of the problem: adapters are not granular experts for individual queries, but specialized modules designed to solve specific tasks.
[…]
Concretely, LoRAuter first constructs a database of representative tasks using small, publicly available validation sets. For each task, it identifies the most suitable adapter from the pool using efficient search strategies, such as Successive Halving (Jamieson & Talwalkar, 2016), which substantially reduces evaluation cost. At inference time, incoming queries are embedded and matched against task representations, allowing LoRAuter to retrieve the top-$K$ relevant tasks and their associated adapters. These adapters are then composed using a weighted output-space fusion mechanism that reflects input–task similarity, enabling effective generalization to both in-domain and out-of-domain queries.
We implement LoRAuter and evaluate our framework on a mixed-task benchmark comprising diverse tasks and $48$ adapters Zhao et al. (2024). Our results demonstrate that LoRAuter achieves 101.2101.2% of the performance of an Oracle task-aligned adapter in in-domain settings, effectively outperforming the upper bound of selecting the perfect adapter for a query. In out-of-domain scenarios, where the model must generalize to unseen tasks, LoRAuter outperforms the strongest baseline, LoraRetriever Zhao et al. (2024), by $5.2$ percentage points. We then scale our system to a pool of 1500+ adapters fetched from the wild, and show that LoRAuter achieves competitive performance to the original $48$ adapters on the same mixed-task benchmark. Thus, LoRAuter can extract useful routing signals from large, noisy adapter collections, offering a scalable and robust solution for open-ended LoRA serving.
Our main contributions are:
  • Training-free, black-box routing: We introduce a LoRA routing framework that requires neither adapter training data nor router model training, enabling black-box use of large and heterogeneous adapter pools.
  • Efficiency: LoRAuter scales efficiently with the number of tasks, which is typically much smaller than the number of available adapters. We also employ a search strategy (Successive Halving) to quickly identify optimal adapters for representative tasks, reducing computational overhead by over $2\times$ compared to full adapter evaluation.
  • Extensive evaluation: We conduct comprehensive evaluations across large adapter pools, multiple model sizes, and ablation settings, demonstrating the effectiveness of LoRAuter.
2 Background and problem description
Table 1: A qualitative comparison of recently proposed model routing approaches. $N,T$ and $L$ refer to the total adapters in the pool, number of representative tasks, and number of layers in the model. We typically have $T<N$.
Method Training data not required Training free Adapter Selection Overhead
MoLE  $_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{ [ICLR24]}}$ Wu et al. (2024) $\surd$ $\times$ $O(NL)$
LoraHub  $_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{ [COLM24]}}$ Huang et al. (2024) $\surd$ $\times$ $O(1)$
AdapterSoup  $_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{ [FindingsEACL23]}}$ Chronopoulou et al. (2023) $\times$ $\surd$ $O(1)$
LoraRetriever  $_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{ [FindingsACL24]}}$ Zhao et al. (2024) $\times$ $\surd$ $O(N)$
Arrow $_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{[ICML24]}}$ Ostapenko et al. (2024) $\surd$ $\surd$ $O(NL)$
LoRAuter (Ours) $\surd$ $\surd$ $O(T)$
[…]
LoRA (Hu et al., 2022) is a parameter-efficient fine-tuning (PEFT) technique that enables adapting LLMs to excel on downstream tasks with minimal additional parameters. Instead of updating all model weights, LoRA trains a pair of low-rank matrices for selected weight layers, while keeping the original weights frozen. Formally, for a weight matrix $W\in\mathbb{R}^{m\times d}$, LoRA learns two smaller matrices $A\in\mathbb{R}^{r\times d}$ and $B\in\mathbb{R}^{m\times r}$ such that the effective parameter update is given by $\Delta W=BA$, where $r\ll\min(m,d)$ denotes the adaptation rank. During inference, the forward computation becomes $h^{\prime}=(W+\Delta W)x=Wx+BAx$, allowing efficient adaptation while keeping the number of new parameters small. LoRA achieves competitive performance compared to full fine-tuning but substantially reduces storage and computation costs.
2.2 Composing LoRA adapters
[…]
$h^{\prime}=Wx+(\lambda\Delta W_{1}+(1-\lambda)\Delta W_{2})x,$ (1)
where $\lambda\in[0,1]$ controls the relative contribution of each adapter. However, this approach requires manual tuning of $\lambda$ for each task. More advanced approaches, such as LoraHub (Huang et al., 2024), extend this idea by learning adaptive fusion weights rather than relying on manually tuned coefficients. Given a set of LoRA adapters, the fused representation is computed as:
$h^{\prime}=Wx+\left(\sum_{i}w_{i}B_{i}\right)\left(\sum_{i}w_{i}A_{i}\right)x,$ (2)
where $w_{i}$ are learnable weights optimized in a few-shot manner. Although LoraHub effectively automates the fusion process, the learned weights still remain task-specific, requiring separate optimization for each new task and incurring additional computational overhead.
2.3 Problem formulation
We formalize the problem of LoRA adapter routing. We consider a pool of $N$ LoRA adapters, $\Phi=\{\phi_{1},\phi_{2},\ldots,\phi_{N}\}$, available for routing, where each adapter $\phi_{i}=(B_{i},A_{i})$. In practice, an adapter is not made of a single pair of matrices, but of a pair of matrices for every layer of the model that benefits from the adapter Hu et al. (2022). To simplify notation, we consider these adapter layers to be implicit.
[…]
The goal of routing is to efficiently select and compose adapters from the pool to respond to a given user query, generalizing to unseen tasks or domains. In this sense, we seek an open-ended multi-LoRA serving framework, capable of dynamically adapting to heterogeneous requests as new LoRA adapters are added to the pool, extending the vision outlined by Zhao et al. (2024). Under this scenario, given a query $x$ without its task label, the serving process can be expressed as: $y=F(\mathcal{R}_{x},x,W)$ where $W$ refers to the weights of base LLM, $\mathcal{R}_{x}$ is the output of the selection or retrieval mechanism for query $x$ and $F$ is the composition function. For instance, $\mathcal{R}_{x}$ can correspond to the top-$K$ LoRA adapters retrieved for answering $x$ and $F$ can correspond to one of the aforementioned fusion strategies.
2.4 Shortcomings of existing approaches
Table 1 summarizes existing LoRA adapter routing approaches along key dimensions, including whether they (i) require access to adapter training data, (ii) involve additional training for routing or composition, and (iii) incur inference-time overhead that scales with the number of adapters $(N)$ and model layers $(L)$. Our objective is a routing mechanism that is training-free, operates in a black-box setting without access to adapter training data, and remains computationally efficient as the adapter pool grows.
Many prior methods violate one or more of these desiderata. Mixture-of-Experts (MoE)-style approaches such as MoLE Wu et al. (2024) rely on learned gating networks trained with task-specific data, which must be retrained for each task and limits applicability in open-ended settings. Similarly, LoraHub Huang et al. (2024) learns adapter composition weights using validation data, requiring task-specific supervision at deployment time.
Other approaches avoid training but still depend on adapter training data. AdapterSoup Chronopoulou et al. (2023) and LoraRetriever Zhao et al. (2024) embed adapters using training data and perform routing by matching queries to these representations. While effective, such methods are unsuitable for public or proprietary adapter pools where training data is unavailable. In contrast, spectral routing methods such as Arrow Ostapenko et al. (2024) eliminate both training data and learned routing components by operating directly on adapter parameters, but incur higher adapter selection overheads due to layer-wise routing.
Beyond data and training requirements, existing methods also differ substantially in inference-time selection overhead. Methods like LoraHub and AdapterSoup produce a single composite adapter per task and incur constant selection cost, while MoLE and Arrow perform routing at every layer, leading to $O(NL)$ complexity. LoraRetriever avoids layer-wise routing but still compares each query against all adapters, resulting in $O(N)$ complexity.
In contrast, LoRAuter operates at the level of tasks rather than individual adapters. By routing queries to task representations, LoRAuter reduces selection complexity to $O(T)$, where $T$ denotes the number of representative tasks (Section 3.1) and typically satisfies $T<N$. This design enables scalable, training-free routing that remains efficient as the adapter pool grows.
3 Design of LoRAuter
Figure 1: The LoRAuter framework.
We now present the design of LoRAuter. The central idea of LoRAuter is to perform LoRA adapter selection and composition at the level of tasks, rather than directly reasoning over individual adapters. To this end, LoRAuter is organized into four core components: (1) a database of relevant tasks, (2) task-to-adapter pairing, (3) query-based task retrieval, and (4) adapter composition. The overall workflow of LoRAuter is illustrated in Figure 1.
3.1 Database of relevant tasks ()
LoRAuter begins by constructing a representative task database $\mathcal{T}^{rep}=\{t_{1},t_{2},\ldots,t_{T}\}$ from publicly available task queries. This task database serves as a repository of $T$ diverse tasks representing the range of query types the system may encounter. Each task $t_{i}$ is accompanied by a small validation set $D_{i}$ containing labeled samples, which are lightweight and readily available in practice today. Note that $T$ can be typically much lower than $N$, the total number of adapters in the pool.
3.2 Task to adapter pairing (➌, ➍, ➎)
[…]
Since both the number of tasks and available adapters can be substantial, and most adapters are only effective for a limited subset of tasks, exhaustive evaluation quickly becomes computationally prohibitive. To address this, we employ tournament-based search strategies, such as Successive Halving (Jamieson & Talwalkar, 2016) (SH), which iteratively eliminate underperforming adapters while performing additional evaluations on more promising candidates. This efficiently identifies the best adapter for each task while substantially reducing computational overhead. Furthermore, SH scales to hundreds of adapters by parallelizing evaluation rounds across GPUs with minimal coordination.
Alternative methods, such as LoraHub (Huang et al., 2024), can also be used to determine optimal adapters by learning fusion weights. However, these techniques exhibit limited scalability, as the training process for adapter fusion is typically more computationally expensive than adapter evaluation. Notably, our framework can easily incorporate such training-based methods by treating the fused adapter produced by LoraHub (or similar approaches) as an additional candidate within the adapter pool. If the fused adapter surpasses the current best adapter for a task, it becomes the new selected adapter. We leave the integration and evaluation of these methods as directions for future work.
3.3 Query-based task retrieval (➏, ➐, ➑, ➒)
LoRAuter identifies the most relevant task for a given query using task representations. These representations are constructed in a one-time offline manner by employing a sentence-embedding model $E$. Specifically, the model $E$ encodes a subset of the tasks’ validation queries to obtain the task representation as follows.
For each task $t_{i}$, we construct a fixed embedding by encoding and aggregating a small subset of its validation queries. Let $\{v_{j}\}_{j=1}^{m}\subset D_{i}$ be a randomly sampled subset with $m\ll|D_{i}|$. Given an embedding instruction $I$ and a validation query $v$, the encoder $E$ produces an embedding by encoding the concatenation $I\oplus v$. The task representation $\mathbf{e}_{i}$ is then computed as:
$\mathbf{e}_{i}=\frac{1}{m}\sum_{j=1}^{m}E(I\oplus v_{j}).$
Similar to Zhao et al. (2024), we set $I$ to the following instruction: “Represent the sentence for similar task retrieval”, allowing the embedding to capture task similarity.
At inference time, we similarly encode an arriving query $x$ to obtain an embedding $\mathbf{e}_{x}$. This embedding is then compared against the precomputed task embeddings $\{\mathbf{e}_{i}\}_{i=1}^{T}$ using the cosine similarity metric:
$s_{i}=\frac{\mathbf{e}_{x}\cdot\mathbf{e}_{i}}{\|\mathbf{e}_{x}\|\,\|\mathbf{e}_{i}\|},$ (3)
where $s_{i}$ denotes the similarity between the input and the task $t_{i}$. We then retrieve the set of top-$K$ most similar tasks based on their scores $s_{i}$: $\mathcal{T}_{x}=\{t_{i}\mid s_{i}\text{ is in top-}K\}$. To obtain a probabilistic interpretation of the input-task affinities, we normalize these scores using the softmax function with a temperature parameter $\tau$. The resulting probability $p_{i}$ reflects the likelihood that the input belongs to task $t_{i}$. The sentence encoder used in this paper is trained using the Supervised Contrastive (SupCon) loss (Khosla et al., 2020), which minimizes a temperature-scaled cross-entropy objective over cosine similarities. Thus, applying a softmax function to cosine similarities yields a probabilistic interpretation consistent with the encoder’s training objective.
3.4 Adapter composition ()
After retrieving the top-$K$ tasks $\mathcal{T}_{x}$ for a given query $x$ along with corresponding probabilities $\{p_{i}\}_{i=1}^{K}$, we use the precomputed task-to-adapter mapping $\mathcal{M}$ to identify the adapter $\phi\in\Phi$ associated with each task $t_{i}$. We define $\mathcal{R}_{x}=\{(p_{i},\mathcal{M}[t_{i}])\mid t_{i}\in\mathcal{T}_{x}\}$ as the output of selection step, containing relevant adapters coupled with their matching likelihoods. The goal is now to compose these $K$ adapters to produce the final output, leveraging their complementary strengths.
While prior methods Zhao et al. (2024); Chronopoulou et al. (2023) treat all retrieved tasks as equally relevant by using uniform averaging for composition, we use an input-aware fusion mechanism that dynamically adjusts the contribution of each adapter based on the input-task similarity. Specifically, the fused adapter outputs $h^{\prime}$ are obtained through a weighted combination of the top-$K$ adapters in the output space:
$h^{\prime}=Wx+\sum_{(w_{i},\phi_{i})\in\mathcal{R}_{x}}w_{i}\,B_{i}A_{i}x,$ (4)
where $\sum_{i=1}^{K}w_{i}=1$ where the weights $w_{i}$ correspond to probabilities $p_{i}$. This formulation adaptively integrates information from task-relevant adapters, allowing the model to emphasize those most aligned with the input while minimizing interference from unrelated tasks. Furthermore, composition in the output space yields better performance than composition in the parameter space Zhao et al. (2024), as LoRA adapters often exhibit limited alignment, which hinders effective parameter merging Stoica et al. (2025).
4 Evaluation
We next evaluate the effectiveness of the LoRAuter framework using representative benchmarks and against four state-of-the-art routing baselines. Our evaluation proceeds as follows: we first report main performance results under non-OOD, OOD, and semi-OOD settings (see Figure 6, Appendix A), and study robustness when scaling to large adapter pools (§4.2); we then conduct two ablation studies to assess the contributions of retrieval and composition (§4.3); we examine performance when task labels are replaced with data-driven pseudo-tasks obtained via clustering (§4.4); and finally, we evaluate the efficiency of adapter selection using Successive Halving (§4.5).
4.1 Experimental setup
Dataset. We adopt a curated subset of 48 individual tasks from Flanv2 (Wei et al., 2022), covering a broad range of Natural Language Understanding (NLU) and Natural Language Generation (NLG) settings. We note that the same set of tasks is used in Zhao et al. (2024), enabling a direct comparison between LoRAuter and LoraRetriever.
Adapters. We use LLaMA 2-7B,13B (Touvron et al., 2023b) as the base model, following prior work and leveraging its large ecosystem of publicly available LoRA adapters. Each task from the above dataset is treated independently and used to train a dedicated LoRA adapter, resulting in a pool of 48 task-specific adapters. Each LoRA adapter is trained using the Alpaca instruction-following format (Taori et al., 2023). The rank $r$ and scaling hyperparameter $\alpha$ are fixed to 6 and 12, respectively, matching the configuration in the LoraRetriever baseline. Following standard LoRA practice, the learned low-rank weight update is scaled by $\alpha/r$. To further ensure comparability, we also adopt the same text encoder used in Zhao et al. (2024) for generating task embeddings and input representations. Lastly, we also assess LoRAuter on a public adapter pool fetched from HuggingFace, comprising 1500+ adapters.
[…]
Baselines. We compare LoRAuter against the following approaches: LoraHub (Huang et al., 2024), a universal fusion method that learns global mixture weights across; LoraRetriever (Zhao et al., 2024), a retrieval-based method that ranks adapters using input to training data similarity scores and performs fusion over the top-$K$ most similar adapters using uniform weights (mixture); ARROW (Ostapenko et al., 2024), a spectral routing approach that applies SVD to adapter representations and selects adapters based on spectral affinity with the input; and SpectR (Spectral Routing) (Fleshman & Durme, 2025), an eigenspace-based spectral routing variant that projects inputs and adapters into aligned eigenspaces for similarity-based selection. As an oracle, we report the performance achieved when directly using the adapter trained on the same task. Note that the oracle is fixed, and remains independent of the OOD or non-OOD setting.
Metrics. Following Wei et al. (2022) and Zhao et al. (2024), task performance is measured using task-specific metrics: Exact Match (EM) accuracy for NLU tasks, BLEU for translation tasks, and ROUGE-1, ROUGE-2, and ROUGE-L for structure-to-text generation tasks. To aggregate performance across tasks with different metrics, we normalize each score by dividing it by the oracle performance for the corresponding task. We then average the normalized scores across tasks and report the resulting metric.
Figure 2: Average normalized performance of different routing methods on non-OOD and OOD tasks for both models.
4.2 Main results
OOD and non-OOD settings. Figure 2 presents the results for both the models (Llama2-7B & Llama2-13B) and settings. The corresponding detailed numerical results are included in Table 6 (Appendix B.3). Our proposed LoRAuter achieves the strongest overall performance across each configuration. Under non-OOD setting and Llama2-7B, it attains an average normalized score of 101.2%, slightly surpassing the oracle (100%). This shows that uncertainty-aware fusion of complementary adapters from related tasks can sometimes even outperform the adapter trained for a given task. Under the OOD setting and Llama2-7B, LoRAuter reaches 88.4%, outperforming LoraRetriever, the best baseline by +5.2 points. The same pattern holds for Llama2-13B. LoRAuter achieves 98.8% on non-OOD evaluation, again closely matching the oracle, and 86.8% in the OOD setting. Spectral routing methods (Arrow and SpectR) achieve relatively lower performance as they directly route based on parameter values, which may not carry sufficient routing signal. As with the 7B model, LoRAuter outperforms baselines in each setting, demonstrating that the advantages of uncertainty-aware composition persist even on larger models.
Semi-OOD setting. We introduce a semi-OOD setting that lies between the non-OOD and OOD regimes and reflects realistic deployment scenarios. In this setting, the adapter trained on the target task is unavailable, but a validation set from the target task is accessible for routing. Figure 3 presents the results, with detailed numerical values reported in Table 6 (Appendix B.3). With access to validation data, LoRAuter can more accurately identify relevant adapters using task representations. For LoraRetriever, we continue to report OOD performance, as it does not leverage validation data and relies solely on training data. We find that LoRAuter improves to 92.7% in the semi-OOD setting from 88.4% in the OOD setting on Llama2-7B. Similarly, it improves to 91.1% from 86.8% on Llama2-13B. On both models, this corresponds to gains exceeding 4 points and remains substantially higher than LoraRetriever.
Figure 3: Performance in the semi-OOD setting where validation data is available but the task-specific adapter is removed.
Scaling the adapter pool.
We study the robustness of LoRAuter under large and heterogeneous adapter pool fetched “from the wild”. We collect all adapters available for Llama2-7B with rank $\leq 64$ from HuggingFace, resulting in a pool of 1567 adapters. We run LoRAuter on this pool which now does not contain the 48 well curated adapters but only adapters fetched from the wild. In the OOD setting, LoRAuter impressively achieves an average normalized performance of $85.7\%$ compared to $88.4\%$ obtained using the 48 adapter pool. Similarly, in the semi-OOD setting, LoRAuter achieves $89.6\%$ compared to $92.7\%$ on the well curated pool. This result demonstrates the potential of LoRAuter to extract useful routing signals from large, noisy adapter collections with no information about their training data whatsoever. To the best of our knowledge, this is the first work to evaluate LoRA adapter routing using such large collections of publicly available adapters. In this setting, the baseline methods cannot be meaningfully evaluated, as they either require access to adapter training data or remain computationally expensive to scale to 1500+ adapters. The detailed numerical results are included in Table 7 (Appendix B.3); for completeness, we also report results for a combined adapter pool that augments the HuggingFace adapters with the original 48 curated adapters.
Table 2: Performance on OOD tasks. Both the retrieval and composition components of LoRAuter (LA) individually bring performance improvements over the baseline LoraRetriever (LR).
Model Retrieval Composition Norm. Performance
non-OOD OOD
Llama2-7B LR LR 92.9 83.2
LR LA 98.6 85.7
LA LR 96.7 89.9
LA LA 101.2 88.4
Llama2-13B LR LR 95.6 85.9
LR LA 99.1 86.2
LA LR 96.8 86.3
LA LA 98.8 86.8
Table 3: The effect of composition for both LoraRetriever (LR) and LoRAuter (LA) on OOD tasks.
Model Method w/o Composition (K=1) with Composition (K=3)
Llama2-7B LR 78.4 83.2
LA 81.8 88.4
Llama2-13B LR 76.7 85.9
LA 78.3 86.8
4.3 Ablation Studies
We conduct two ablation studies to assess: (i) the individual effects of the retrieval and composition components of LoRAuter (LA), and (ii) the benefit of composition compared to selecting only the single best adapter. For the first ablation, we replace the retrieval and composition components of our baseline LoraRetriever (LR) with the corresponding components of LA. Table 2 shows that just using the weighted composition scheme of LA improves LR from 92.9% to 98.6% on Llama2-7B (non-OOD). Similarly, using only the retrieval component of LA improves LR from 92.9% to 96.7% on the same setup. The best performance of 101.2% is achieved when both the components are used. We observe a similar trend for the OOD setup with Llama2-13B. On the other hand, for LLama2-7B in OOD setting and LLama2-13B in non-OOD setting, peak performance is achieved when replacing only one component of LR with LA. Nevertheless, the performance of using both LA components remains very close and emerges as globally best in two out of the four cases. Corresponding to the second ablation, Table 3 reports the performance of using the single best adapter $(K=1)$ as retrieved by both LR and LA. As shown in the table, this settings yields fairly low performance compared to the $K=3$ case, demonstrating the importance of composing multiple relevant adapters for a given query in the OOD scenario.
4.4 LoRAuter with pseudo tasks
We study how LoRAuter behaves when ground-truth task labels are not available for the publicly collected queries for constructing the representative task set. Instead of relying on the original 48 task labels, we apply K-Means to the validation data representations and partition the data into $K$ pseudo task clusters. We repeat the evaluation for different values of $K$ while keeping the total number of validation samples fixed. In this setting, the routing mechanism must rely purely on the structure of the data, without access to true task annotations.
Figure 4: Normalized performance of LoRAuter with varying number of clusters $K$.
Figure 4 shows that the normalized performance remains within a stable range and peaks at an intermediate number of clusters. Too few clusters produce overly coarse partitions that fail to capture meaningful variation, while too many yield small, noisy clusters that provide unreliable signals for adapter selection. The optimal choice of $K$ varies across regimes. For OOD inputs, fewer (larger) clusters perform better, encouraging routing toward robust adapters that generalize across broader data distributions. In contrast, non-OOD inputs benefit from finer-grained clusters that enable selection of more specialized adapters.
Importantly, even without access to ground truth task labels, LoRAuter achieves performance comparable to routing with task labels by appropriately tuning the number of K-Means clusters. At the best-performing configurations, the method attains up to 101.5% (7B) and 99.3% (13B) normalized performance for non-OOD inputs, and 88.0% (7B) and 87.1% (13B) for OOD inputs.
4.5 Searching the best adapter
We now compare two strategies for identifying the best adapter for a given task under a fixed compute budget (the number of adapter calls). We consider: (i) Uniform Selection: the total evaluation budget is evenly divided across all adapters, and each adapter is evaluated on the same number of validation samples before selecting the highest-scoring one; (ii) Successive Halving: adapters are progressively filtered based on performance estimates, allocating more evaluation budget to promising candidates while respecting the same total budget. For each budget, we conduct 100 independent runs, where each run randomly samples a subset from the validation datasets. Figure 5 shows the normalized performance of both methods across varying evaluation budgets. Successive Halving (SH) reduces the evaluation budget to reach very close to peak performance by more than 2 times as compared to Uniform Selection. Thus, LoRAuter scales efficiently to larger adapter pools by avoiding evaluation of all adapters on all samples.
Figure 5: Normalized performance of uniform selection vs. Successive Halving (SH) across evaluation budgets for Llama2-7B.
5 Related work
Mixture-of-Experts.
MoE (Jacobs et al., 1991; Jordan & Jacobs, 1993; Shazeer et al., 2017) architectures partition a model into multiple experts and use a learned gating mechanism to sparsely activate experts for each input or token. This enables expert specialization while scaling model capacity without a proportional increase in inference cost. MoE systems are typically trained end-to-end, with expert parameters and routing functions jointly optimized and fixed after training. Recent work extends this paradigm to PEFT by treating LoRA adapters as experts and learning routers to select among them (Muqeeth et al., 2024; Wu et al., 2024). While effective, these approaches require training additional routing parameters and assume a static set of adapters during router training. In contrast, LoRAuter is training-free.
Model routing. Model-routing methods aim to reduce inference cost by dynamically selecting an appropriate model for each input while maintaining response quality. Prior work (Wang et al., 2025b) distinguishes between non-predictive and predictive routing, where the prior relies on intermediate model outputs to guide routing decisions such as via early exiting (Chen et al., 2024; Aggarwal et al., 2024), while the latter estimates response quality before inference to directly select a model (Ding et al., 2024; Wang et al., 2025b). While most approaches assume small, static model pools, recent work such as ICL-Router (Wang et al., 2025a) extends routing to more dynamic settings using learned model capability embeddings. However, these methods treat models as indivisible units, selecting a single model per query, which is misaligned with modern PEFT where a single base model may expose thousands of specialized LoRA adapters and where composing multiple adapters can be beneficial. In contrast, our framework performs predictive routing at the adapter level using task representations and supports input-aware adapter fusion, enabling scalable routing over large adapter collections.
Task representations. Prior work has explored representing entire tasks, rather than individual inputs, as compact vector representations, commonly referred to as task representations. Such representations have been used to characterize task similarity and predict transferability across tasks (Achille et al., 2019; Zamir et al., 2018). More recently, MoE-based routing architectures have incorporated task-related signals to guide expert selection (Liang et al., 2025). In contrast, existing approaches to LoRA adapter routing primarily operate at the level of individual adapters, constructing adapter representations from either their parameters or training data, and do not expose tasks as explicit, reusable routing primitives. While LoraRetriever (Zhao et al., 2024) implicitly leverages task-related information for input-aware adapter retrieval, tasks are not represented as first-class entities and therefore cannot be independently associated with or shared across adapters.
6 Conclusions
We introduced LoRAuter, a training-free framework for scalable LoRA adapter routing that operates at the level of tasks rather than individual adapters. By combining task representations, efficient task–adapter pairing, and input-aware adapter composition, LoRAuter enables mixed-task inference without access to adapter training data and with overhead that scales in the number of tasks. Empirically, LoRAuter matches oracle task-aligned performance in-domain and achieves state-of-the-art results out-of-domain. Overall, LoRAuter provides a practical foundation for leveraging large, dynamic adapter pools for LoRA routing.
Impact statement
This work advances Machine Learning by enabling scalable, training-free routing and composition of large collections of LoRA adapters. By selectively reusing and fusing task-relevant adapters, our framework can improve the performance of smaller base language models, reducing the need to rely on larger and more memory-intensive LLMs. This is particularly beneficial for users operating under GPU memory constraints, where using large models may be impractical or prohibitively expensive. More broadly, our framework is designed to benefit the open-source community by allowing practitioners to reuse and combine thousands of publicly available adapters for new tasks without requiring training data or additional fine-tuning. This can substantially lower computational and financial barriers, making high-quality, task-adaptive LLM capabilities more accessible to a wider range of users, including academic researchers, small labs, and independent developers.
Our method inherits the strengths and weaknesses of the underlying base models and adapters, including potential biases, privacy risks, or harmful behaviors. By making it easier to rapidly reuse and combine public adapters, it could also unintentionally enable misuse in sensitive domains. We therefore encourage responsible deployment, including careful vetting of adapters, appropriate safety filters, and consideration of ethical and regulatory requirements.
Appendix A Evaluation details
This section provides additional evaluation details for all methods considered in our study. All baselines are evaluated under an identical experimental setup: each method operates over the same pool of 48 LoRA adapters with fixed rank $r=6$ and scaling $\alpha=12$, and is assessed on the same mixed-task evaluation set. All experiments are conducted using both LLaMA2-7B and LLaMA2-13B backbone models, with inference performed using bfloat16 precision. For LoRAuter Fusion, we fuse the $k=3$ adapters and use a softmax temperature of $\tau=0.2$ in all evaluations. The sentence transformer used is https://huggingface.co/Styxxxx/lora_retriever, which is the same as the one found in the LoraRetriever paper.
We evaluate our methods under three different routing regimes, illustrated in Figure 6. In the non-OOD (in-domain) setting, both the validation samples and the LoRA adapter corresponding to the input’s ground-truth task are available during routing. In the Semi-OOD setting, validation samples for the target task remain available, but the adapter trained explicitly on that task is removed from the adapter pool, requiring methods to generalize by selecting adapters trained on related tasks. Finally, in the OOD (out-of-domain) setting, both the task-specific validation samples and the corresponding adapter are excluded, representing the most challenging scenario in which routing must rely entirely on cross-task generalization.
Figure 6: Illustration of the evaluation regimes used in our experiments. non-OOD (left) corresponds to the in-domain setting where both task-specific validation samples and the task-aligned LoRA adapter are available during routing. Semi-OOD (center) removes the aligned adapter while retaining validation samples. OOD (right) excludes both the task-specific validation samples and the corresponding adapter.
A.1 Details of baseline methods
For completeness, we evaluate LoRAuter against several representative training-free or lightweight adapter–routing approaches. All baselines are implemented following the algorithmic descriptions in their respective papers, and where applicable, we reuse the same text encoder and LoRA configurations to ensure a fair comparison. The baselines correspond exactly to those summarized in Section 4.1 of this paper.
LoraHub
Huang et al. (2024). LoraHub is a task-specific adapter–merging method that learns task-specific fusion weights over a set of LoRA parameters via black-box optimization. The method outputs a single fused adapter per task and does not support input-adaptive routing. Following the recommendations in the original paper, we randomly sample 20 adapters from the full pool to constrain the optimization search space. For evaluation, each input is answered using the fused adapter optimized for its corresponding task (i.e., we select the task-specific fused adapter based on the ground-truth task label). LoRAHub was trained using the same validation set as in our method, and we fix the number of optimization iterations to 48, matching the total number of adapter evaluations performed by our approach.
LoraRetriever
Zhao et al. (2024). A retrieve-then-compose method that ranks LoRA adapters using similarity between input and adapter representations, then fuses the top-$k$ adapters uniformly. This method assumes the availability of supervised training data for each task and requires homogeneous adapter training conditions. Our method was evaluated using the exact same set of adapters and the same text encoder as employed in the original LoraRetriever paper.
Arrow
Ostapenko et al. (2024). A spectral routing method that applies SVD to adapter representations and selects experts based on spectral affinity with the input. It is training-free but introduces additional per-token inference overhead proportional to the number of adapters. Following the recommendations of the original paper, we use a temperature of $1$ for converting spectral affinities into routing weights, and we retrieve $k=3$ adapters during routing for consistency with our other baselines.
SpectR
Fleshman & Durme (2025). A recently proposed spectral routing variant that aligns input and adapter representations through eigenspace projections. It avoids training but generally underperforms methods that incorporate task-level context. While the method supports softmax-based weighting of the selected adapters, in our evaluation we follow the configuration used in the original paper and apply uniform weights instead. Furthermore, we retrieve $k=4$ adapters during routing, as suggested in the original paper.
Oracle task-aligned adapter.
For reference, we report an oracle task-aligned baseline, which corresponds to selecting the LoRA adapter that was explicitly trained on the ground-truth task of each input.
Normalized average computation.
To obtain a single aggregate performance metric across heterogeneous tasks and evaluation scales, we compute a normalized average. For structure-to-text generation, we first average Rouge-1, Rouge-2, and Rouge-L to obtain a single score per method. Then, for each task (i.e., each row in the results table), we normalize all method scores using the corresponding value in the Oracle Task-Aligned column as the normalization constant. Finally, the normalized averages reported in Table 2 are computed by averaging these per-task normalized scores across all tasks.
A.2 Dataset details
We adopt the mixed-task evaluation benchmark introduced by Zhao et al. (2024), which is designed to reflect realistic multi-task inference scenarios where user prompts span heterogeneous task types and task labels are not available at inference time. The benchmark comprises a pool of 48 LoRA adapters, each trained independently on a distinct downstream task drawn from a diverse set of natural language understanding (NLU) and natural language generation (NLG) datasets. This setup mirrors practical serving environments in which LoRA adapters are specialized for different tasks.
Task taxonomy.
The 48 tasks covered by the benchmark follow the high-level categorization popularized by the FLAN instruction-tuning framework (Wei et al., 2022). For evaluation and reporting, as illustrated in Figure 7 and reported in Zhao et al. (2024), these 48 tasks are grouped into 10 semantically coherent benchmark tasks. Concretely, the 48 adapters span task categories including natural language inference, reading comprehension, closed-book question answering, commonsense reasoning, sentiment analysis, coreference resolution, paraphrase detection, structure-to-text generation, and machine translation. This taxonomy ensures coverage of both classification-style tasks with discrete label spaces and open-ended generation tasks with free-form outputs.
Figure 7: Overview of the task taxonomy and dataset coverage used in the mixed-task benchmark. Each LoRA adapter is trained on a dataset belonging to a specific task cluster (e.g., natural language inference, reading comprehension, translation), following the taxonomy introduced in FLAN (Wei et al., 2022). At inference time, inputs are provided without task labels, requiring routing methods to infer the underlying task and select or compose suitable adapters. Figure reproduced from LoraRetriever (Zhao et al., 2024).
Instruction formatting.
Following the instruction-based paradigm of Wei et al. (2022), each dataset is verbalized using natural language instructions that describe the task objective (e.g., entailment determination, sentiment classification, or translation). While multiple instruction templates are used during adapter training in the original datasets, our evaluation treats adapters as black-box modules and does not assume access to their training prompts or data. At inference time, inputs are provided as plain user queries without explicit task identifiers, reflecting the mixed-task setting.
Validation and test data.
For each task, a small validation subset is used exclusively for routing or adapter selection, while a disjoint test set is used for final evaluation. Validation sets are intentionally lightweight (typically a few hundred examples per task) to reflect realistic constraints where large-scale labeled data may be unavailable. Test metrics follow standard task-specific evaluation measures (e.g., accuracy, exact match, BLEU, or ROUGE), as reported in the main results.
Overall, this benchmark provides a controlled yet challenging testbed for evaluating scalable LoRA routing methods under both in-domain and distribution-shifted conditions.
Appendix B Additional results
In addition to the primary findings discussed in the main body, this section provides a comprehensive breakdown of the performance of alternative adapter routing strategies. Beyond these comparisons, we also include supplementary adapter evaluation results on the validation set. Finally, we present supplementary results for the Successive Halving (SH) algorithm to further illustrate its efficiency.
B.1 Selection-based methods
In this section, we report results for several selection-based strategies for adapter routing. While these methods achieve strong performance in the non-OOD setting, they underperform relative to weighted fusion-based routing. This gap arises because selection-based approaches are more sensitive to input–task misalignment introduced by the imperfect text encoder used for input-to-task routing.
Moreover, selection-based methods collapse in the Semi-OOD and OOD settings, as they lack a mechanism for aggregating knowledge across multiple related tasks and must rely on a single selected adapter. Table 4 summarizes the performance of the evaluated selection strategies across the non-OOD, Semi-OOD, and OOD settings.
The quantitative results in Table 4 highlight a widening performance gap as the evaluation shifts toward out-of-distribution (OOD) scenarios. In the non-OOD setting, both LoraRetriever and LoRAuter selections maintain high normalized averages (e.g., 99.0% for LoRAuter on Llama2-7B), indicating that the text encoders are generally capable of identifying the correct specialized adapter when the task is known. However, this proficiency diminishes significantly in Semi-OOD and OOD regimes. For instance, the normalized average for LoRAuter Selection drops from 99.0% to 81.8% on Llama2-7B when moving from non-OOD to OOD. Furthermore, while the larger Llama2-13B model exhibits higher absolute scores across most metrics, it follows a similar degradation pattern to the 7B variant, suggesting that increased parameter scale does not inherently resolve the fundamental limitations of hard-selection routing in zero-shot or OOD contexts.
Table 4: Performance of selection-based adapter routing methods across evaluation regimes. Task-Aligned reports the oracle adapter trained on the ground-truth task, while Perf-Aligned selects the single adapter with highest validation performance on the ground-truth task. LoraRetriever Selection and LoRAuter Selection denote input-aware single-adapter selection ($k=1$) using the corresponding routing method.
Llama2-7B
Oracle Non-OOD Semi-OOD OOD
Domain-Metric Task- Aligned Perf- Aligned LoRA- Retriever Selection LoRAuter Selection LoRAuter Selection LoRA- Retriever Selection LoRAuter Selection
struct to text-rouge-1 63.5 63.5 61.3 61.0 49.1 50.1 43.4
struct to text-rouge-2 39.0 39.0 37.0 36.9 25.1 26.6 20.8
struct to text-rouge-l 56.6 56.6 54.5 54.6 43.2 43.9 38.5
commonsense-em 63.0 68.5 55.5 68.5 65.5 46.0 62.0
sentiment-em 90.0 90.5 89.5 90.0 87.5 89.0 90.0
reading comp-em 67.7 67.7 51.7 63.7 53.7 40.3 48.7
closed_book QA-em 44.0 50.5 40.0 45.0 40.0 43.0 41.0
coreference-em 51.0 59.0 50.0 59.0 59.0 46.0 52.0
read.comp.w:com 70.0 70.0 69.0 59.0 36.0 30.0 19.0
paraphrase-em 63.5 69.0 58.0 60.5 56.0 45.5 53.0
nli-em 73.1 72.5 70.0 71.7 67.6 60.6 64.0
translation-bleu 13.1 12.6 12.8 12.6 11.9 12.0 11.7
Normalized Average (%) 100.0 104.4 93.2 99.0 88.4 78.4 81.8
Llama2-13B
Oracle Non-OOD Semi-OOD OOD
Domain-Metric Task- Aligned Perf- Aligned LoRA- Retriever Selection LoRAuter Selection LoRAuter Selection LoRA- Retriever Selection LoRAuter Selection
struct to text-rouge-1 65.0 65.0 62.6 62.6 50.4 49.4 44.1
struct to text-rouge-2 40.4 40.4 38.2 38.2 26.8 25.8 21.9
struct to text-rouge-l 58.3 58.3 56.0 56.0 44.0 42.9 38.9
commonsense-em 69.0 68.5 59.0 66.0 64.0 47.5 65.5
sentiment-em 90.0 90.0 90.5 90.0 90.0 91.0 92.0
reading comp-em 76.3 76.3 60.3 72.3 60.7 48.0 46.0
closed_book QA-em 65.0 64.0 60.0 61.0 58.0 53.0 56.5
coreference-em 75.0 76.0 75.0 76.0 69.0 65.0 55.0
read.comp.w:com 81.0 81.0 80.0 67.0 54.0 33.0 34.0
paraphrase-em 76.5 76.5 68.0 67.0 59.5 52.5 53.5
nli-em 82.0 82.6 78.9 79.6 69.5 70.2 73.5
translation-bleu 12.8 13.2 12.9 13.2 12.9 12.7 12.8
Normalized Average (%) 100.0 100.3 93.8 95.2 85.8 76.7 78.3
Varying the number of task clusters.
In this section, we study how sensitive LoRAuter selection is to the granularity of the task database by varying the number of pseudo-task clusters $K$ used during routing. Following Section 4.4, we remove access to ground-truth task labels and instead apply K-Means on validation-set representations, partitioning the available validation data into $K$ clusters while keeping the total validation budget fixed. LoRAuter then performs retrieval and adapter pairing with respect to these cluster-defined pseudo-tasks. As shown in Figure 6, performance exhibits an inverted-U trend across model sizes and evaluation regimes: using too few clusters produces overly coarse groupings that blur meaningful task structure, whereas using too many clusters yields small, noisy partitions that provide unreliable signals for adapter selection. Moreover, the best-performing $K$ shifts slightly by regime, with the OOD setting favoring fewer (larger) clusters than non-OOD. Under distribution shift LoRAuter benefits from selecting robust adapters that perform well across a broad set of validation samples. In contrast, in the non-OOD regime finer-grained clustering enables LoRAuter to pick more specialized adapters that perform well on a small set of samples.
Under our cluster-based evaluation, the best $K$ yields normalized scores of $98.7\%$ (K=96K{=}96), and $97.4\%$ (K=384K{=}384) for LLaMA2-7B and LLaMA2-13B in the non-OOD setting, while dropping to $85.1\%$, (K=12(K{=}12) and $81.7\%$, (K=12K{=}12) in OOD setting. These trends align with the results in 4 for LoRAuter Selection, which report $99.0\%$ (7B) and $95.2\%$ (13B) in non-OOD versus $81.8\%$ (7B) and $78.3\%$ (13B) in OOD.
Figure 8: Normalized performance of LoRAuter selection method ($k=1$) as the number of clusters $K$ varies for both 7B and 13B models. LoRAuter almost reaches the upper bound set by the oracle.
B.2 LoraRetriever weighted fusion
This subsection evaluates the effect of applying our weighted fusion mechanism to the LoraRetriever routing strategy. Table 5 reports results across non-OOD and OOD regimes for both LLaMA2-7B and LLaMA2-13B backbones. Incorporating weighted fusion substantially improves the previous LoraRetriever Mixture method, and achieves performance that is nearly on par with LoRAuter in both the non-OOD and OOD settings. However, unlike LoRAuter, this approach does not exploit additional task samples when they are available. As a result, in the Semi-OOD setting, where limited validation data from the target task can be leveraged for routing, weighted fusion over LoraRetriever cannot adapt its mixture to the new task and therefore significantly underperforms the LoRAuter method.
Comparing normalized averages across the three fusion strategies further clarifies this trend. In non-OOD, LoraRetriever’s uniform mixture reaches $92.9\%$ (LLaMA2-7B) and $95.6\%$ (LLaMA2-13B), while weighted fusion improves to $98.6\%$ and $99.1\%$, approaching LoRAuter Fusion at $101.2\%$ and $98.8\%$, respectively. In OOD, the same ordering holds: the original mixture scores $83.2\%$ (7B) and $85.9\%$ (13B), weighted fusion rises to $85.7\%$ and $86.2\%$, and LoRAuter Fusion remains strongest at $88.4\%$ and $86.8\%$. Overall, weighting reduces errors from imperfect retrieval by upweighting the most relevant retrieved adapters, yielding consistent gains over uniform mixing.
Table 5: Results of applying weighted adapter fusion to LoraRetriever on Non-OOD and OOD tasks for LLaMA2-7B and LLaMA2-13B
Llama2-7B Llama2-13B
Domain-Metric Non-OOD OOD Non-OOD OOD
struct to text-rouge-1 59.5 48.0 60.7 50.3
struct to text-rouge-2 33.6 24.7 36.2 27.1
struct to text-rouge-l 52.7 42.5 53.6 43.9
commonsense-em 60.0 58.0 70.0 67.5
sentiment-em 90.5 91.0 92.0 92.5
reading comp-em 59.3 47.7 70.3 53.0
closed_book QA-em 47.0 46.5 66.0 59.0
coreference-em 58.0 57.0 81.0 62.0
read.comp.w:com 66.0 36.0 79.0 56.0
paraphrase-em 62.5 45.5 71.5 58.0
nli-em 71.0 63.9 80.0 74.9
translation-bleu 13.2 12.2 13.4 13.9
Normalized Average (%) 98.6 85.7 99.1 86.2
B.3 Main results per benchmark task
OOD, non-OOD, semi-OOD settings.
We complement Figure 2 by reporting results broken down by benchmark task. Table 6 and Table 9 present per-benchmark normalized scores, showing that LoRAuter achieves the strongest performance on the majority of tasks. This indicates that the gains observed in Figure 2 are not driven by a single task category, but are broadly consistent across diverse benchmarks.
We further observe a similar trend in the semi-OOD setting, where the task-specific adapter is unavailable but validation data for the target task remains accessible. In this regime, LoRAuter continues to outperform baselines, demonstrating its ability to leverage task-level validation signals. Compared to the fully OOD setting, semi-OOD performance improves consistently across benchmarks. Overall, these results confirm that LoRAuter ’s advantages hold at a fine-grained, per-task level and extend robustly across non-OOD, semi-OOD, and OOD scenarios.
Scaling the adapter pool.
The scaling results in Section 4.2 show that LoRAuter maintains performance comparable to baseline routing methods evaluated on the 48-adapter pool, even when operating with adapters collected exclusively “from the wild,” over substantially larger and noisier adapter collections, and without access to well-curated task-specific adapters.
In addition to the HF-only setting discussed in Section 4.2, we also evaluate LoRAuter on a combined adapter pool consisting of the 1,567 publicly available HuggingFace adapters augmented with the original 48 curated task-specific adapters. This setting allows us to assess whether LoRAuter can effectively leverage high-quality task-aligned adapters when they are present, while remaining robust to a large number of adapters. We find that performance in this combined pool closely matches the results obtained with the curated adapters alone, indicating that the presence of a large, noisy adapter pool does not hinder LoRAuter ’s ability to select and compose strong adapters. Detailed per-task results are provided in Table 7 and Table 8.
Table 6: Adapter performance across multiple domains. Non-OOD denotes in-domain tasks, while OOD and Semi-OOD correspond to different out-of-domain evaluation settings. The LoRAuter column represents the proposed input-aware adapter routing method. The Oracle Task-Aligned baseline (shaded in gray) reports performance of the adapter trained on the ground-truth task of each input. Results for the LoraRetriever Mixture are taken directly from the original LoraRetriever paper. All average scores are normalized by the Task-Aligned performance per task.
Llama2-7B
Non-OOD OOD Semi-OOD
Domain-Metric Oracle Task- Aligned LoRA- Retriever Mixture LoRAuter LoRA- Retriever Mixture LoRAuter LoRAuter
struct to text-rouge-1 63.5 55.9 58.8 50.4 48.5 51.4
struct to text-rouge-2 39.0 30.0 33.6 26.9 25.3 27.3
struct to text-rouge-l 56.6 49.5 51.8 44.0 43.1 45.3
commonsense-em 63.0 61.5 68.0 50.0 63.0 65.5
sentiment-em 90.0 89.5 90.5 90.5 91.0 91.0
reading comp-em 67.7 51.3 62.3 47.3 47.3 51.7
closed_book QA-em 44.0 45.0 50.5 48.5 47.5 46.0
coreference-em 51.0 63.0 63.0 49.0 54.0 65.0
read.comp.w:com 70.0 46.0 67.0 40.0 39.0 42.0
paraphrase-em 63.5 56.5 60.0 46.0 53.5 57.5
nli-em 73.1 67.9 70.3 56.5 64.0 69.3
translation-bleu 13.1 12.8 12.6 12.2 12.9 11.8
Normalized Average (%) 100.0 92.9 101.2 83.2 88.4 92.7
Llama2-13B
Non-OOD OOD Semi-OOD
Domain-Metric Oracle Task- Aligned LoRA- Retriever Mixture LoRAuter LoRA- Retriever Mixture LoRAuter LoRAuter
struct to text-rouge-1 65.0 57.7 61.1 52.1 50.7 53.3
struct to text-rouge-2 40.4 32.6 36.5 28.1 27.2 29.3
struct to text-rouge-l 58.3 50.8 53.9 45.4 43.8 47.0
commonsense-em 69.0 64.0 70.5 60.5 67.0 69.5
sentiment-em 90.0 91.5 91.5 91.5 92.5 91.0
reading comp-em 76.3 60.3 71.3 51.3 52.3 60.7
closed_book QA-em 65.0 63.0 63.0 61.0 59.5 60.0
coreference-em 75.0 76.0 77.0 64.0 66.0 68.0
read.comp.w:commonsense-em 81.0 78.0 79.0 58.0 56.0 63.0
paraphrase-em 76.5 71.0 71.0 55.5 59.0 66.0
nli-em 82.0 78.1 79.5 75.7 74.7 75.7
translation-bleu 12.8 14.6 14.3 14.1 13.9 13.9
Normalized Average (%) 100.0 95.6 98.8 85.9 86.8 91.0
Table 7: Per-domain results for LoRAuter on Llama2-7B across Non-OOD, Semi-OOD and OOD settings for two different adapter pools: only 48 well curated adapters (48), only adapters from the wild (HF), both well curated and from the wild adapters (HF+48)
HF+48 HF
Domain-Metric Non-OOD Semi-OOD OOD Semi-OOD OOD
struct to text-rouge-1 58.8 51.2 49.6 48.7 36.2
struct to text-rouge-2 33.6 27.1 25.6 24.2 18.1
struct to text-rouge-l 52.0 44.7 43.9 43.5 32.0
commonsense-em 68.0 65.5 64.0 65.5 63.5
sentiment-em 90.5 90.5 91.5 89.0 88.5
reading comp-em 61.7 53.3 48.3 51.0 49.7
closed_book QA-em 48.5 46.0 47.0 41.5 37.0
coreference-em 60.0 59.0 54.0 62.0 64.0
read.comp.w:com 68.0 44.0 46.0 34.0 37.0
paraphrase-em 61.0 57.0 52.5 38.5 52.5
nli-em 69.9 68.2 68.0 69.6 70.3
translation-bleu 12.4 10.9 11.8 11.5 5.2
Table 8: Normalized average performance results for LoRAuter on Llama2-7B across Non-OOD, Semi-OOD and OOD settings for three adapter pools: only 48 well curated adapters (48), only adapters from the wild (HF), both well curated and from the wild adapters (HF+48)
Setting Non-OOD Semi-OOD OOD
48 101.2 92.7 88.4
HF / 89.6 85.7
HF + 48 100.2 91.4 89.6
Table 9: Baseline adapter performance across multiple domains. Non-OOD denotes in-domain tasks, while OOD correspond to out-of-domain evaluation setting.
Llama2-7B
Non-OOD OOD
Domain-Metric LoRAHub ARROW SpectR LoRAHub ARROW SpectR
struct to text-rouge-1 55.0 39.5 44.2 44.8 40.1 44.2
struct to text-rouge-2 31.9 19.6 22.5 23.1 20.1 22.5
struct to text-rouge-l 48.0 35.5 39.5 39.8 35.6 39.5
commonsense-em 44.5 59.0 47.0 42.5 54.0 47.0
sentiment-em 61.0 82.5 69.5 63.5 86.0 69.5
reading comp-em 44.0 51.7 43.0 45.0 50.7 43.0
closed_book QA-em 34.5 46.5 37.0 34.0 47.0 37.0
coreference-em 43.0 60.0 57.0 46.0 59.0 57.0
read.comp.w:com 15.0 31.0 25.0 14.0 27.0 25.0
paraphrase-em 35.0 45.0 42.5 38.0 49.5 42.5
nli-em 49.2 56.2 53.5 50.6 59.9 53.5
translation-bleu 11.9 11.6 11.6 11.7 10.8 11.6
Normalized Average (%) 68.6 82.5 74.2 67.8 82.0 66.3
Llama2-13B
Non-OOD OOD
Domain-Metric LoRAHub ARROW SpectR LoRAHub ARROW SpectR
struct to text-rouge-1 57.1 45.3 48.0 44.4 46.1 48.0
struct to text-rouge-2 34.8 23.6 25.0 22.9 23.8 25.0
struct to text-rouge-l 50.9 40.1 42.0 39.0 39.9 42.0
commonsense-em 55.5 68.5 59.5 56.0 64.0 59.5
sentiment-em 43.0 90.5 75.0 41.5 91.0 75.0
reading comp-em 47.3 55.7 51.7 48.0 53.0 51.7
closed_book QA-em 59.0 64.5 54.5 59.5 62.5 54.5
coreference-em 45.0 61.0 53.0 48.0 61.0 53.0
read.comp.w:com 30.0 59.0 39.0 30.0 52.0 39.0
paraphrase-em 38.5 52.5 44.5 36.5 54.5 44.5
nli-em 48.5 70.2 59.8 49.6 70.7 59.8
translation-bleu 13.7 14.1 14.0 13.5 13.1 14.0
Normalized Average (%) 68.2 85.7 75.1 66.1 83.2 69.9
B.4 Adapter performance on the validation set
In this subsection, we analyze the normalized performance of each adapter evaluated on the validation set. Figure 9 visualizes this performance matrix for both LLaMA2-7B and LLaMA2-13B backbone models. Each row corresponds to a single evaluation task, while each column represents a single LoRA adapter trained for a specific task.
To facilitate comparison across tasks with heterogeneous metrics and score ranges, we apply row-wise normalization. Specifically, for each task (row), we subtract the minimum observed score across all adapters and divide by the range between the maximum and minimum scores for that task, resulting in values normalized to the interval $[0,1]$. This normalization emphasizes relative adapter performance within each task while removing scale effects induced by differing evaluation metrics.
The figure reveals a pronounced diagonal trend, indicating that adapters trained on a given task generally achieve strong performance when evaluated on that same task. However, this diagonal dominance is not absolute: in several cases, adapters trained on related tasks outperform the task-specific adapter.
In addition to task-specific peaks, we observe consistent blocks of elevated performance among adapters belonging to the same task group. This effect is especially visible for the structure-to-text and machine translation tasks in the bottom-right of the figure, where adapters form well-defined blocks with strong intra-group generalization, reflecting both their shared structure and their use of different evaluation metrics compared to classification-style tasks.
Figure 9: Performance matrix of adapters across tasks evaluated on the validation data. Each row corresponds to one task, while each column represents a single LoRA adapter trained for a specific task.
B.5 Successive halving
To efficiently evaluate the effectiveness of Successive Halving (SH) for adapter selection, we reuse pre-computed adapter outputs obtained from up to 200 validation samples per task. We exclude tasks with fewer than 200 validation examples to ensure that every run can draw sufficiently large subsets while keeping the evaluation protocol consistent across tasks. To compare performance across heterogeneous metrics and scales, we normalize each task by the same Oracle Task-Aligned domain normalization constants as before, and report the resulting normalized average across tasks.
We estimate the SH and uniform evaluation curves using 100 independent runs per evaluation budget. In each run, we randomly sample subsets of varying sizes from the original pool of 200 validation samples, perform adapter selection under the corresponding budget, and then evaluate the selected adapter on the task’s held-out test set.
Figure 10 shows that for both LLaMA2-7B and LLaMA2-13B, Successive Halving consistently achieves comparable normalized test performance to uniform evaluation while requiring substantially fewer validation evaluations. In particular, SH reduces the average number of evaluations per adapter by more than $\times 2$, demonstrating that tournament-style elimination can significantly lower selection cost without sacrificing downstream performance.
Figure 10: Normalized performance of Successive Halving over uniform selection with respect to their allocated compute budget, for the LLama-7B model (left) and 13B model (right). Standard deviation is reported over the randomness of the validation set ordering with 100 samples.
Pseudo-code.
Algorithm 1 summarizes the Successive Halving (SH) procedure used for adapter selection under a fixed evaluation budget (measured in adapter-sample evaluations). Compared to the standard formulation, we (i) use a fixed per-adapter evaluation budget for the first $k$ rounds (warmup), and (ii) decouple the elimination rate from the budget schedule by using separate constants for retention ($\eta$) and per-round budget growth ($\gamma$). Starting from an initial candidate set, SH evaluates each adapter on a validation subset, retains the top fraction, and repeats with an increased budget allocation until a single adapter remains.
Algorithm 1 Successive Halving
Input: adapters $\mathcal{A}$, validation set $\mathcal{V}$, base samples $m$, keep ratio $\eta\in(0,1)$, budget growth $\gamma>1$, rounds $R$, warmup rounds $k$
Output: selected adapter $a^{\star}$
$\mathcal{S}\leftarrow\mathcal{A}$
for $r=0$ to R1R-1 do
  if $r<k$ then
   $m_{r}\leftarrow m$ // fixed budget during warmup
  else
   $m_{r}\leftarrow\lceil m\cdot\gamma^{\,r-k+1}\rceil$ // grow budget after warmup
  end if
  $m_{r}\leftarrow\min(m_{r},\ |\mathcal{V}|)$
  Sample $S_{r}\subseteq\mathcal{V}$ with $|S_{r}|=m_{r}$
  for all $a\in\mathcal{S}$ do
   $s[a]\leftarrow\mathrm{Score}(a,S_{r})$
  end for
  $q\leftarrow\max\!\left(1,\left\lceil\eta\cdot|\mathcal{S}|\right\rceil\right)$
  $\mathcal{S}\leftarrow$ top-$q$ adapters in $\mathcal{S}$ by $s[\cdot]$
  if $|\mathcal{S}|=1$ then
   break
  end if
end for
$a^{\star}\leftarrow$ the single remaining adapter in $\mathcal{S}$
B.6 Choosing $K$ for Input-Aware Adapter Fusion
In Table 10, we evaluate how the number of tasks selected for each input ($K$) affects LoRAuter performance with Llama2-7B in the in-distribution setting. We vary $K$ from $1$ (selecting a single task, i.e., using one adapter without any fusion) to $5$ (fusing the top-$5$ task-specific adapters per input). The best performance is achieved at $K=2$, with $K=3$ performing essentially identically. We therefore select $K=3$ throughout the rest of the paper to remain consistent with prior fusion-based baselines.
Table 10: Ablation with different values of $K$. Based on these results, we select $K=3$, which achieves performance close to the best result ($K=2$) while remaining consistent with baselines which also use $K=3$.
Value of $K$ 1 2 3 4 5
struct to text-rouge-1 60.9 59.5 58.4 58.5 58.9
struct to text-rouge-2 36.6 34.2 33.1 33.2 33.6
struct to text-rouge-l 54.4 52.9 51.3 51.8 51.7
commonsense-em 68.0 69.5 67.5 67.0 68.5
sentiment-em 91.0 90.5 90.5 90.5 90.5
reading comp-em 64.0 63.7 62.7 61.0 61.3
closed_book QA-em 45.0 48.5 49.5 49.0 51.5
coreference-em 58.0 62.0 66.0 63.0 65.0
read.comp.w:commonsense-em 59.0 69.0 67.0 62.0 61.0
paraphrase-em 61.5 62.0 60.5 60.5 61.5
nli-em 70.9 69.5 70.1 70.5 70.5
translation-bleu 12.5 12.8 12.7 12.7 12.7
Normalized Average (%) 98.9 101.8 101.6 99.9 101.2
Appendix C Full results
In Tables 11 and 12, we report detailed per-task results for LoRAuter across each of the 48 tasks in the evaluation dataset. Results are reported for both LLaMA2-7B and LLaMA2-13B backbone models and cover all evaluation regimes considered in the study.
In Tables 13 and 14, we report per-task results for the LoRAHub, ARROW, and SpectR baselines on both LLaMA2-7B and LLaMA2-13B, evaluated under the Non-OOD, Semi-OOD, and OOD regimes.
In Table 15, we report per-task results for LoRAuter on Llama2-7B using 1,615 publicly available HuggingFace adapters with rank $\leq 64$, evaluated under Non-OOD, Semi-OOD, and OOD regimes.
In Tables 16 and 17, we present detailed per-task results for selection-based routing methods on both LLaMA2-7B and LLaMA2-13B across the same set of 48 evaluation tasks. These tables include results for oracle task-aligned selection, performance-aligned selection, and input-aware selection strategies based on both LoraRetriever and LoRAuter.
In Table 18, we present detailed per-task results for the LoraRetriever weighted-mixture routing method on both LLaMA2-7B and LLaMA2-13B. All results follow the same evaluation protocol, enabling direct comparison across methods and evaluation settings.
Table 11: Per-task results for fusion-based adapter routing methods on LLaMA2-7B across Non-OOD, Semi-OOD, and OOD evaluation regimes. Results for the LoraRetriever Mixture baseline are directly copied from the original LoraRetriever paper.
Non-OOD Semi-OOD OOD
Domain-Metric Oracle Task- Aligned LoRA- Retriever Mixture LoRAuter Fusion LoRAuter Fusion LoRA- Retriever Mixture LoRAuter Fusion
Struct to Text
web_nlg_en-rouge-1 69.6 57.8 63.4 52.3 53.9 49.9
web_nlg_en-rouge-2 48.0 33.5 41.9 29.0 29.4 28.0
web_nlg_en-rouge-l 62.8 52.3 56.3 48.1 49.6 45.4
dart-rouge-1 71.7 63.2 60.1 60.3 60.0 57.4
dart-rouge-2 49.4 36.6 34.4 35.7 35.4 33.6
dart-rouge-l 64.9 56.3 53.1 54.1 52.4 51.8
e2e_nlg-rouge-1 66.1 66.0 64.1 63.1 58.7 55.6
e2e_nlg-rouge-2 39.6 38.8 37.6 36.0 32.1 30.4
e2e_nlg-rouge-l 56.4 56.9 54.0 52.9 49.0 48.6
common_gen-rouge-1 46.5 36.5 47.8 30.1 29.0 31.0
common_gen-rouge-2 18.9 11.1 20.6 8.4 8.6 9.2
common_gen-rouge-l 42.2 32.7 43.6 26.0 24.8 26.7
Translation
para_crawl_enes-bleu 24.0 22.8 23.9 19.5 22.1 23.6
wmt16_translate_tren-bleu 3.7 3.7 2.0 2.0 2.6 2.3
wmt16_translate_deen-bleu 17.8 11.0 19.5 18.5 10.8 20.4
wmt16_translate_ruen-bleu 11.5 18.8 11.9 11.3 18.7 12.0
wmt16_translate_fien-bleu 6.8 7.3 7.0 5.9 7.8 6.6
wmt16_translate_roen-bleu 14.6 13.1 12.4 13.2 12.2 13.8
wmt14_enfr-bleu 16.0 17.8 16.8 16.8 18.0 17.9
wmt16_translate_csen-bleu 10.5 8.3 7.3 7.2 5.8 6.5
COMMONSENSE
story_cloze-em 72.0 84.0 96.0 96.0 58.0 88.0
piqa-em 44.0 38.0 44.0 38.0 34.0 42.0
copa-em 86.0 80.0 82.0 80.0 68.0 78.0
hellaswag-em 50.0 44.0 50.0 48.0 40.0 44.0
sentiment
sst2-em 98.0 96.0 98.0 98.0 94.0 94.0
yelp_polarity_reviews-em 98.0 98.0 98.0 98.0 98.0 98.0
imdb_reviews-em 96.0 96.0 98.0 98.0 96.0 98.0
sentiment140-em 68.0 68.0 68.0 70.0 74.0 74.0
READING Comp.
multirc-em 68.0 48.0 64.0 56.0 44.0 44.0
squad_v2-em 62.0 22.0 54.0 20.0 16.0 14.0
squad_v1-em 68.0 62.0 72.0 68.0 68.0 64.0
openbookqa-em 84.0 78.0 80.0 76.0 66.0 74.0
bool_q-em 84.0 80.0 82.0 74.0 76.0 74.0
drop-em 40.0 18.0 22.0 16.0 14.0 14.0
CLOSE-BOOK QA
natural_questions-em 18.0 16.0 14.0 12.0 10.0 12.0
arc_easy-em 48.0 66.0 78.0 76.0 82.0 82.0
arc_challenge-em 46.0 50.0 56.0 50.0 46.0 48.0
trivia_qa-em 64.0 48.0 54.0 46.0 56.0 48.0
COREFERENCE
definite_pronoun_resolution-em 52.0 68.0 70.0 72.0 56.0 58.0
wsc-em 50.0 58.0 56.0 58.0 42.0 50.0
READ. COMP. W/ COM
cosmos_qa-em 70.0 50.0 72.0 52.0 46.0 48.0
record-em 70.0 42.0 62.0 32.0 34.0 30.0
PARAPHRASE
paws_wiki-em 90.0 56.0 60.0 66.0 46.0 54.0
glue_qqp-em 70.0 80.0 78.0 78.0 58.0 74.0
glue_mrpc-em 60.0 60.0 66.0 66.0 58.0 62.0
stsb-em 34.0 30.0 36.0 20.0 20.0 24.0
NLI
cb-em 88.9 86.7 88.9 86.7 66.7 57.8
wnli-em 72.0 60.0 56.0 58.0 54.0 50.0
anli_r1-em 52.0 40.0 54.0 50.0 42.0 48.0
anli_r2-em 46.0 46.0 46.0 44.0 46.0 48.0
anli_r3-em 46.0 44.0 46.0 46.0 50.0 48.0
mnli_matched-em 88.0 80.0 86.0 86.0 88.0 86.0
mnli_mismatched-em 92.0 88.0 94.0 92.0 90.0 94.0
snli-em 96.0 90.0 84.0 84.0 92.0 82.0
qnli-em 94.0 74.0 70.0 68.0 38.0 66.0
rte-em 56.0 70.0 78.0 78.0 76.0 60.0
Table 12: Per-task results for fusion-based adapter routing methods on LLaMA2-13B across Non-OOD, Semi-OOD, and OOD evaluation regimes. Results for the LoraRetriever Mixture baseline are directly copied from the original LoraRetriever paper.
Non-OOD Semi-OOD OOD
Domain-Metric Oracle Task- Aligned LoRA- Retriever Mixture LoRAuter Fusion LoRAuter Fusion LoRA- Retriever Mixture LoRAuter Fusion
Struct to Text
web_nlg_en-rouge-1 72.4 59.7 68.9 57.4 53.7 52.4
web_nlg_en-rouge-2 52.1 35.6 46.9 31.6 29.1 28.9
web_nlg_en-rouge-l 66.1 55.1 61.0 50.8 49.0 46.6
dart-rouge-1 72.9 62.6 64.3 59.0 60.6 59.2
dart-rouge-2 52.9 38.9 40.4 35.4 37.3 35.6
dart-rouge-l 66.7 55.0 56.4 53.1 53.4 51.7
e2e_nlg-rouge-1 66.5 66.7 65.1 64.4 63.9 61.6
e2e_nlg-rouge-2 38.8 39.0 37.4 36.7 36.4 33.9
e2e_nlg-rouge-l 56.4 56.2 55.5 53.4 53.7 51.6
common_gen-rouge-1 48.4 41.7 46.3 32.5 30.0 29.7
common_gen-rouge-2 17.9 17.0 21.2 13.6 9.5 10.6
common_gen-rouge-l 44.1 36.7 42.8 30.5 25.3 25.4
Translation
para_crawl_enes-bleu 25.0 27.4 25.9 24.3 25.8 23.6
wmt16_translate_tren-bleu 2.0 3.4 3.7 3.6 2.7 2.8
wmt16_translate_deen-bleu 19.8 10.5 21.3 21.7 12.0 20.9
wmt16_translate_ruen-bleu 11.3 20.2 12.4 11.9 20.2 11.9
wmt16_translate_fien-bleu 6.8 8.7 9.3 8.2 7.8 8.6
wmt16_translate_roen-bleu 14.2 17.4 11.7 11.2 13.3 13.7
wmt14_enfr-bleu 16.5 18.7 20.5 20.5 20.9 20.9
wmt16_translate_csen-bleu 6.9 10.3 9.5 9.7 10.1 9.0
COMMONSENSE
story_cloze-em 94.0 80.0 98.0 96.0 76.0 96.0
piqa-em 48.0 46.0 50.0 48.0 46.0 48.0
copa-em 76.0 78.0 80.0 82.0 76.0 80.0
hellaswag-em 58.0 52.0 54.0 52.0 44.0 44.0
sentiment
sst2-em 98.0 98.0 98.0 96.0 100.0 100.0
yelp_polarity_reviews-em 98.0 98.0 98.0 98.0 98.0 100.0
imdb_reviews-em 96.0 98.0 98.0 98.0 98.0 98.0
sentiment140-em 68.0 72.0 72.0 72.0 70.0 72.0
READING Comp.
multirc-em 88.0 66.0 78.0 66.0 44.0 40.0
squad_v2-em 68.0 60.0 68.0 38.0 60.0 18.0
squad_v1-em 74.0 34.0 70.0 70.0 24.0 64.0
openbookqa-em 86.0 88.0 88.0 78.0 82.0 78.0
bool_q-em 84.0 78.0 84.0 86.0 78.0 86.0
drop-em 58.0 36.0 40.0 26.0 20.0 28.0
CLOSE-BOOK QA
natural_questions-em 30.0 24.0 26.0 18.0 16.0 18.0
arc_easy-em 94.0 94.0 94.0 90.0 94.0 94.0
arc_challenge-em 68.0 68.0 70.0 70.0 70.0 70.0
trivia_qa-em 68.0 66.0 62.0 62.0 64.0 56.0
COREFERENCE
definite_pronoun_resolution-em 88.0 88.0 90.0 76.0 72.0 74.0
wsc-em 62.0 64.0 64.0 60.0 56.0 58.0
READ. COMP. W/ COM
cosmos_qa-em 84.0 82.0 82.0 80.0 70.0 60.0
record-em 78.0 74.0 76.0 46.0 46.0 18.0
PARAPHRASE
paws_wiki-em 92.0 88.0 88.0 78.0 74.0 64.0
glue_qqp-em 86.0 84.0 82.0 82.0 62.0 82.0
glue_mrpc-em 84.0 78.0 72.0 68.0 64.0 64.0
stsb-em 44.0 34.0 42.0 36.0 22.0 26.0
NLI
cb-em 95.6 93.3 93.3 91.1 88.9 86.7
wnli-em 72.0 76.0 74.0 68.0 66.0 64.0
anli_r1-em 70.0 64.0 66.0 64.0 68.0 68.0
anli_r2-em 66.0 60.0 62.0 54.0 60.0 56.0
anli_r3-em 68.0 62.0 66.0 62.0 60.0 60.0
mnli_matched-em 86.0 86.0 86.0 86.0 88.0 84.0
mnli_mismatched-em 90.0 94.0 92.0 92.0 100.0 92.0
snli-em 90.0 90.0 92.0 90.0 92.0 92.0
qnli-em 94.0 74.0 84.0 76.0 58.0 74.0
rte-em 88.0 82.0 80.0 74.0 76.0 70.0
Table 13: Performance of baseline routing methods on Llama2-7B .
Non-OOD OOD
Domain-Metric LoRAHub ARROW SpectR LoRAHub ARROW SpectR
Struct to Text
web_nlg_en-rouge-1 67.2 30.8 41.6 46.6 30.6 33.6
web_nlg_en-rouge-2 44.4 15.8 22.0 24.2 15.5 16.9
web_nlg_en-rouge-l 59.1 29.4 39.0 42.6 28.2 31.6
dart-rouge-1 61.2 49.4 53.6 60.4 50.8 44.0
dart-rouge-2 36.7 26.3 30.3 36.0 29.4 24.1
dart-rouge-l 54.7 44.4 48.3 53.7 46.2 39.5
e2e_nlg-rouge-1 65.1 49.7 53.5 47.2 52.1 49.2
e2e_nlg-rouge-2 38.6 27.2 29.3 25.3 27.7 26.7
e2e_nlg-rouge-l 55.1 42.4 45.4 40.8 43.5 42.0
common_gen-rouge-1 26.3 28.2 28.1 24.9 26.9 25.7
common_gen-rouge-2 7.8 9.0 8.3 6.8 7.8 6.2
common_gen-rouge-l 23.0 25.7 25.4 22.2 24.7 24.0
Translation
para_crawl_enes-bleu 23.6 21.6 23.4 22.6 19.5 19.6
wmt16_translate_tren-bleu 3.5 2.7 2.4 2.4 3.8 2.9
wmt16_translate_deen-bleu 16.9 17.6 16.6 16.4 17.5 12.6
wmt16_translate_ruen-bleu 10.1 9.4 9.1 9.7 6.8 6.9
wmt16_translate_fien-bleu 7.3 7.2 6.6 8.4 5.9 5.9
wmt16_translate_roen-bleu 10.6 10.3 11.3 10.9 9.3 8.4
wmt14_enfr-bleu 16.6 17.1 17.0 16.9 16.8 15.0
wmt16_translate_csen-bleu 6.4 7.0 6.4 6.5 6.9 6.8
COMMONSENSE
story_cloze-em 62.0 84.0 70.0 64.0 70.0 74.0
piqa-em 30.0 38.0 34.0 28.0 40.0 32.0
copa-em 62.0 74.0 66.0 66.0 70.0 70.0
hellaswag-em 24.0 40.0 18.0 12.0 36.0 12.0
sentiment
sst2-em 66.0 82.0 70.0 72.0 90.0 50.0
yelp_polarity_reviews-em 72.0 94.0 82.0 70.0 96.0 76.0
imdb_reviews-em 50.0 92.0 70.0 52.0 96.0 62.0
sentiment140-em 56.0 62.0 56.0 60.0 62.0 50.0
READING Comp.
multirc-em 42.0 46.0 38.0 44.0 50.0 42.0
squad_v2-em 24.0 24.0 22.0 26.0 22.0 28.0
squad_v1-em 66.0 74.0 62.0 66.0 72.0 48.0
openbookqa-em 52.0 66.0 56.0 54.0 66.0 58.0
bool_q-em 72.0 82.0 70.0 70.0 80.0 66.0
drop-em 8.0 18.0 10.0 10.0 14.0 4.0
CLOSE-BOOK QA
natural_questions-em 14.0 20.0 12.0 12.0 20.0 10.0
arc_easy-em 48.0 64.0 56.0 48.0 66.0 52.0
arc_challenge-em 30.0 44.0 34.0 30.0 44.0 32.0
trivia_qa-em 46.0 58.0 46.0 46.0 58.0 36.0
COREFERENCE
definite_pronoun_resolution-em 50.0 68.0 62.0 56.0 66.0 58.0
wsc-em 36.0 52.0 52.0 36.0 52.0 52.0
READ. COMP. W/ COM
cosmos_qa-em 24.0 32.0 38.0 22.0 34.0 10.0
record-em 6.0 30.0 12.0 6.0 20.0 6.0
PARAPHRASE
paws_wiki-em 46.0 48.0 42.0 44.0 60.0 48.0
glue_qqp-em 20.0 52.0 52.0 32.0 70.0 58.0
glue_mrpc-em 62.0 60.0 62.0 62.0 52.0 60.0
stsb-em 12.0 20.0 14.0 14.0 16.0 8.0
NLI
cb-em 44.4 75.6 71.1 57.8 73.3 80.0
wnli-em 44.0 52.0 54.0 46.0 48.0 46.0
anli_r1-em 36.0 32.0 36.0 34.0 38.0 34.0
anli_r2-em 34.0 32.0 36.0 34.0 40.0 28.0
anli_r3-em 44.0 38.0 40.0 44.0 40.0 36.0
mnli_matched-em 56.0 60.0 54.0 56.0 70.0 44.0
mnli_mismatched-em 76.0 88.0 72.0 74.0 86.0 60.0
snli-em 64.0 70.0 62.0 64.0 74.0 58.0
qnli-em 46.0 58.0 52.0 46.0 68.0 56.0
rte-em 48.0 56.0 58.0 50.0 62.0 58.0
Table 14: Performance of baseline routing methods on Llama2-13B.
Non-OOD OOD
Domain-Metric LoRAHub ARROW SpectR LoRAHub ARROW SpectR
Struct to Text
web_nlg_en-rouge-1 71.9 39.8 47.3 42.5 37.7 41.4
web_nlg_en-rouge-2 51.9 20.1 23.5 21.9 17.6 20.1
web_nlg_en-rouge-l 66.1 36.3 42.3 38.9 34.0 38.2
dart-rouge-1 63.0 56.5 56.5 63.0 57.3 45.6
dart-rouge-2 37.8 34.1 33.6 38.0 33.2 27.2
dart-rouge-l 55.9 51.5 51.1 55.7 49.6 40.2
e2e_nlg-rouge-1 66.9 55.9 57.6 49.3 56.6 57.1
e2e_nlg-rouge-2 39.1 29.5 31.0 23.8 30.6 30.3
e2e_nlg-rouge-l 56.6 45.4 46.5 40.0 45.2 45.7
common_gen-rouge-1 26.7 29.0 30.4 22.9 32.8 29.3
common_gen-rouge-2 10.3 10.6 12.0 8.1 13.7 11.4
common_gen-rouge-l 24.9 27.4 28.2 21.4 30.9 27.3
Translation
para_crawl_enes-bleu 22.7 24.5 26.9 22.8 23.2 23.4
wmt16_translate_tren-bleu 4.0 3.9 3.3 3.3 3.6 3.3
wmt16_translate_deen-bleu 18.7 18.9 19.8 18.5 16.2 18.9
wmt16_translate_ruen-bleu 12.2 11.9 11.2 13.0 10.0 10.5
wmt16_translate_fien-bleu 9.9 8.9 8.2 9.9 8.2 6.9
wmt16_translate_roen-bleu 12.1 13.5 12.5 11.8 13.3 10.5
wmt14_enfr-bleu 19.3 19.1 19.6 18.3 18.8 18.9
wmt16_translate_csen-bleu 11.1 11.8 10.7 10.7 11.5 11.5
COMMONSENSE
story_cloze-em 78.0 96.0 92.0 78.0 94.0 90.0
piqa-em 42.0 48.0 44.0 42.0 42.0 44.0
copa-em 68.0 86.0 72.0 68.0 78.0 70.0
hellaswag-em 34.0 44.0 30.0 36.0 42.0 22.0
sentiment
sst2-em 32.0 98.0 78.0 34.0 100.0 70.0
yelp_polarity_reviews-em 46.0 98.0 88.0 46.0 98.0 86.0
imdb_reviews-em 48.0 96.0 68.0 42.0 98.0 62.0
sentiment140-em 46.0 70.0 66.0 44.0 68.0 72.0
READING Comp.
multirc-em 38.0 48.0 54.0 34.0 42.0 46.0
squad_v2-em 34.0 26.0 26.0 34.0 24.0 30.0
squad_v1-em 60.0 68.0 62.0 60.0 68.0 48.0
openbookqa-em 68.0 76.0 68.0 68.0 72.0 70.0
bool_q-em 64.0 88.0 82.0 72.0 84.0 78.0
drop-em 20.0 28.0 18.0 20.0 28.0 8.0
CLOSE-BOOK QA
natural_questions-em 16.0 24.0 14.0 14.0 22.0 16.0
arc_easy-em 94.0 94.0 88.0 94.0 94.0 80.0
arc_challenge-em 58.0 70.0 54.0 62.0 64.0 52.0
trivia_qa-em 68.0 70.0 62.0 68.0 70.0 58.0
COREFERENCE
definite_pronoun_resolution-em 54.0 76.0 64.0 62.0 70.0 70.0
wsc-em 36.0 46.0 42.0 34.0 52.0 42.0
READ. COMP. W/ COM
cosmos_qa-em 42.0 82.0 64.0 42.0 70.0 52.0
record-em 18.0 36.0 14.0 18.0 34.0 14.0
PARAPHRASE
paws_wiki-em 42.0 62.0 52.0 42.0 68.0 44.0
glue_qqp-em 42.0 58.0 56.0 34.0 68.0 42.0
glue_mrpc-em 56.0 64.0 58.0 56.0 62.0 60.0
stsb-em 14.0 26.0 12.0 14.0 20.0 4.0
NLI
cb-em 48.9 82.2 75.6 44.4 86.7 57.8
wnli-em 44.0 54.0 68.0 44.0 60.0 44.0
anli_r1-em 46.0 48.0 36.0 46.0 48.0 34.0
anli_r2-em 30.0 50.0 34.0 32.0 42.0 34.0
anli_r3-em 30.0 50.0 32.0 30.0 50.0 30.0
mnli_matched-em 60.0 92.0 86.0 64.0 92.0 86.0
mnli_mismatched-em 64.0 94.0 74.0 68.0 94.0 78.0
snli-em 52.0 94.0 64.0 52.0 94.0 58.0
qnli-em 50.0 70.0 58.0 54.0 72.0 58.0
rte-em 60.0 68.0 70.0 62.0 68.0 72.0
Table 15: Per-task results of LoRAuter for Llama2-7B with two large adapter pools.
HF+48 HF
Domain-Metric Non-OOD Semi-OOD OOD Semi-OOD OOD
Struct to Text
web_nlg_en-rouge-1 63.6 52.7 50.4 32.3 46.8
web_nlg_en-rouge-2 42.4 29.4 27.6 17.4 23.7
web_nlg_en-rouge-l 57.5 48.2 45.6 29.7 43.3
dart-rouge-1 61.0 61.0 58.9 41.3 57.2
dart-rouge-2 35.7 35.6 33.7 20.9 31.2
dart-rouge-l 53.7 54.0 52.7 38.2 52.2
e2e_nlg-rouge-1 63.7 63.2 58.0 48.4 58.2
e2e_nlg-rouge-2 37.2 36.5 33.0 26.7 31.6
e2e_nlg-rouge-l 53.6 52.3 50.6 39.4 50.8
common_gen-rouge-1 47.0 28.1 31.0 22.9 32.5
common_gen-rouge-2 19.1 7.1 8.1 7.4 10.5
common_gen-rouge-l 43.0 24.3 26.6 20.6 27.6
Translation
para_crawl_enes-bleu 23.5 16.1 22.3 8.3 23.4
wmt16_translate_tren-bleu 2.0 2.0 3.2 1.0 2.3
wmt16_translate_deen-bleu 19.4 17.8 19.9 9.1 16.0
wmt16_translate_ruen-bleu 11.6 8.1 8.7 2.0 10.5
wmt16_translate_fien-bleu 6.2 7.2 6.9 2.5 5.0
wmt16_translate_roen-bleu 13.3 14.1 11.0 4.7 10.9
wmt14_enfr-bleu 16.0 16.2 17.9 9.8 17.1
wmt16_translate_csen-bleu 6.9 5.6 4.3 4.5 6.7
COMMONSENSE
story_cloze-em 96.0 90.0 88.0 94.0 92.0
piqa-em 44.0 40.0 40.0 38.0 38.0
copa-em 82.0 82.0 82.0 80.0 80.0
hellaswag-em 50.0 50.0 46.0 42.0 52.0
sentiment
sst2-em 98.0 98.0 98.0 94.0 92.0
yelp_polarity_reviews-em 98.0 98.0 98.0 98.0 96.0
imdb_reviews-em 98.0 98.0 100.0 98.0 100.0
sentiment140-em 68.0 68.0 70.0 64.0 68.0
READING Comp.
multirc-em 62.0 56.0 42.0 54.0 50.0
squad_v2-em 54.0 20.0 18.0 32.0 32.0
squad_v1-em 74.0 72.0 66.0 54.0 66.0
openbookqa-em 80.0 78.0 76.0 66.0 66.0
bool_q-em 78.0 80.0 72.0 82.0 76.0
drop-em 22.0 14.0 16.0 10.0 16.0
CLOSE-BOOK QA
natural_questions-em 16.0 12.0 12.0 4.0 10.0
arc_easy-em 78.0 78.0 84.0 70.0 70.0
arc_challenge-em 50.0 48.0 46.0 38.0 38.0
trivia_qa-em 50.0 46.0 46.0 36.0 48.0
COREFERENCE
definite_pronoun_resolution-em 62.0 64.0 62.0 70.0 66.0
wsc-em 58.0 54.0 46.0 58.0 58.0
READ. COMP. W/ COM
cosmos_qa-em 72.0 54.0 62.0 52.0 46.0
record-em 64.0 34.0 30.0 22.0 22.0
PARAPHRASE
paws_wiki-em 60.0 66.0 52.0 66.0 54.0
glue_qqp-em 80.0 78.0 76.0 82.0 52.0
glue_mrpc-em 66.0 66.0 62.0 50.0 34.0
stsb-em 38.0 18.0 20.0 12.0 14.0
NLI
cb-em 88.9 84.4 84.4 86.7 84.4
wnli-em 52.0 52.0 54.0 62.0 64.0
anli_r1-em 54.0 50.0 46.0 52.0 52.0
anli_r2-em 44.0 40.0 40.0 40.0 38.0
anli_r3-em 44.0 54.0 56.0 54.0 48.0
mnli_matched-em 88.0 86.0 86.0 82.0 90.0
mnli_mismatched-em 94.0 94.0 92.0 98.0 98.0
snli-em 84.0 84.0 84.0 90.0 88.0
qnli-em 66.0 72.0 72.0 74.0 70.0
rte-em 84.0 66.0 66.0 64.0 64.0
Table 16: Per-task results for selection-based adapter routing methods on LLaMA2-7B, including oracle, performance-aligned, and input-aware selection. Results for LoraRetriever-based selection are directly copied from the original LoraRetriever paper.
Llama2-7B
Oracle Non-OOD Semi-OOD OOD
Domain-Metric Task- Aligned Perf- Aligned LoRA- Retriever Selection LoRAuter Selection LoRAuter Selection LoRA- Retriever Selection LoRAuter Selection
Struct to Text
web_nlg_en-rouge-1 69.6 69.6 67.0 66.2 54.5 53.9 51.4
web_nlg_en-rouge-2 48.0 48.0 44.5 44.8 30.0 30.0 27.8
web_nlg_en-rouge-l 62.8 62.8 60.9 59.9 49.2 49.1 46.7
dart-rouge-1 71.7 71.7 67.9 67.6 60.2 58.4 51.3
dart-rouge-2 49.4 49.4 45.8 46.3 35.2 34.9 27.9
dart-rouge-l 64.9 64.9 61.1 61.8 53.5 52.4 46.3
e2e_nlg-rouge-1 66.1 66.1 65.8 63.9 52.2 59.3 41.5
e2e_nlg-rouge-2 39.6 39.6 39.4 37.4 27.6 34.1 19.6
e2e_nlg-rouge-l 56.4 56.4 55.7 54.2 46.2 50.2 36.5
common_gen-rouge-1 46.5 46.5 44.7 46.2 29.3 29.0 29.5
common_gen-rouge-2 18.9 18.9 18.3 18.9 7.5 7.3 7.8
common_gen-rouge-l 42.2 42.2 40.5 42.4 24.0 24.0 24.4
Translation
para_crawl_enes-bleu 24.0 24.0 24.2 23.8 19.3 20.3 19.6
wmt16_translate_tren-bleu 3.7 2.0 3.1 2.0 2.0 2.6 1.9
wmt16_translate_deen-bleu 17.8 17.8 10.4 17.8 16.9 9.8 16.3
wmt16_translate_ruen-bleu 11.5 11.5 18.7 11.5 11.9 20.3 11.3
wmt16_translate_fien-bleu 6.8 8.0 6.5 8.0 8.0 7.0 7.4
wmt16_translate_roen-bleu 14.6 13.5 14.0 13.5 13.5 12.3 11.9
wmt14_enfr-bleu 16.0 17.0 16.1 17.0 17.0 16.9 18.0
wmt16_translate_csen-bleu 10.5 6.9 9.4 6.9 6.9 7.0 7.0
COMMONSENSE
story_cloze-em 72.0 96.0 62.0 96.0 96.0 42.0 84.0
piqa-em 44.0 44.0 46.0 44.0 34.0 32.0 36.0
copa-em 86.0 84.0 74.0 84.0 84.0 68.0 78.0
hellaswag-em 50.0 50.0 40.0 50.0 48.0 42.0 50.0
sentiment
sst2-em 98.0 98.0 98.0 98.0 96.0 96.0 96.0
yelp_polarity_reviews-em 98.0 98.0 94.0 98.0 98.0 94.0 100.0
imdb_reviews-em 96.0 98.0 96.0 96.0 94.0 96.0 96.0
sentiment140-em 68.0 68.0 70.0 68.0 62.0 70.0 68.0
READING Comp.
multirc-em 68.0 68.0 52.0 68.0 66.0 38.0 62.0
squad_v2-em 62.0 62.0 56.0 62.0 20.0 12.0 14.0
squad_v1-em 68.0 68.0 66.0 68.0 72.0 68.0 62.0
openbookqa-em 84.0 84.0 68.0 76.0 66.0 58.0 76.0
bool_q-em 84.0 84.0 60.0 80.0 78.0 60.0 72.0
drop-em 40.0 40.0 8.0 28.0 20.0 6.0 6.0
CLOSE-BOOK QA
natural_questions-em 18.0 18.0 16.0 16.0 10.0 10.0 10.0
arc_easy-em 48.0 74.0 56.0 74.0 66.0 70.0 74.0
arc_challenge-em 46.0 46.0 42.0 46.0 40.0 46.0 46.0
trivia_qa-em 64.0 64.0 46.0 44.0 44.0 46.0 34.0
COREFERENCE
definite_pronoun_resolution-em 52.0 64.0 50.0 64.0 64.0 50.0 50.0
wsc-em 50.0 54.0 50.0 54.0 54.0 42.0 54.0
READ. COMP. W/ COM
cosmos_qa-em 70.0 70.0 68.0 64.0 50.0 34.0 18.0
record-em 70.0 70.0 70.0 54.0 22.0 26.0 20.0
PARAPHRASE
paws_wiki-em 90.0 90.0 64.0 60.0 52.0 40.0 50.0
glue_qqp-em 70.0 82.0 74.0 82.0 82.0 68.0 78.0
glue_mrpc-em 60.0 70.0 58.0 66.0 72.0 58.0 58.0
stsb-em 34.0 34.0 36.0 34.0 18.0 16.0 26.0
NLI
cb-em 88.9 88.9 80.0 86.7 75.6 62.2 57.8
wnli-em 72.0 64.0 68.0 64.0 64.0 46.0 46.0
anli_r1-em 52.0 52.0 50.0 48.0 46.0 50.0 44.0
anli_r2-em 46.0 38.0 46.0 42.0 38.0 46.0 46.0
anli_r3-em 46.0 38.0 42.0 44.0 44.0 38.0 54.0
mnli_matched-em 88.0 88.0 84.0 88.0 88.0 88.0 88.0
mnli_mismatched-em 92.0 92.0 90.0 92.0 84.0 94.0 94.0
snli-em 96.0 84.0 84.0 84.0 82.0 84.0 82.0
qnli-em 94.0 94.0 94.0 88.0 68.0 26.0 64.0
rte-em 56.0 86.0 62.0 80.0 86.0 72.0 64.0
Table 17: Per-task results for selection-based adapter routing methods on LLaMA2-13B, including oracle, performance-aligned, and input-aware selection. Results for LoraRetriever-based selection are directly copied from the original LoraRetriever paper.
Llama2-7B
Oracle Non-OOD Semi-OOD OOD
Domain-Metric Task- Aligned Perf- Aligned LoRA- Retriever Selection LoRAuter Selection LoRAuter Selection LoRA- Retriever Selection LoRAuter Selection
Struct to Text
web_nlg_en-rouge-1 72.4 72.4 68.5 68.1 56.7 51.9 51.1
web_nlg_en-rouge-2 52.1 52.1 47.5 48.8 33.0 28.6 29.4
web_nlg_en-rouge-l 66.1 66.1 62.4 62.2 50.8 48.3 46.4
dart-rouge-1 72.9 72.9 67.0 69.4 59.7 57.0 51.3
dart-rouge-2 52.9 52.9 45.9 49.5 35.3 33.6 29.2
dart-rouge-l 66.7 66.7 61.2 62.9 53.5 50.0 45.6
e2e_nlg-rouge-1 66.5 66.5 66.1 64.3 57.8 59.2 44.4
e2e_nlg-rouge-2 38.8 38.8 39.3 36.6 31.0 32.8 20.5
e2e_nlg-rouge-l 56.4 56.4 56.4 54.3 45.9 48.9 37.9
common_gen-rouge-1 48.4 48.4 48.9 48.6 27.6 29.3 29.5
common_gen-rouge-2 17.9 17.9 20.3 17.9 7.8 8.2 8.6
common_gen-rouge-l 44.1 44.1 44.1 44.5 25.7 24.4 25.5
Translation
para_crawl_enes-bleu 25.0 25.0 25.4 24.9 22.1 23.1 23.1
wmt16_translate_tren-bleu 2.0 3.2 2.4 3.2 3.2 1.2 2.5
wmt16_translate_deen-bleu 19.8 19.8 11.5 19.8 20.7 10.3 18.0
wmt16_translate_ruen-bleu 11.3 10.9 19.9 10.9 10.9 20.7 11.0
wmt16_translate_fien-bleu 6.8 7.3 6.8 7.3 7.3 5.1 6.6
wmt16_translate_roen-bleu 14.2 10.2 13.9 10.2 10.2 10.9 12.9
wmt14_enfr-bleu 16.5 20.4 17.1 20.4 20.4 18.4 18.8
wmt16_translate_csen-bleu 6.9 8.5 6.2 8.5 8.5 11.6 9.2
COMMONSENSE
story_cloze-em 94.0 94.0 80.0 90.0 90.0 56.0 96.0
piqa-em 48.0 44.0 52.0 44.0 44.0 30.0 40.0
copa-em 76.0 78.0 74.0 78.0 78.0 68.0 78.0
hellaswag-em 58.0 58.0 30.0 52.0 44.0 36.0 48.0
sentiment
sst2-em 98.0 98.0 98.0 98.0 98.0 98.0 100.0
yelp_polarity_reviews-em 98.0 98.0 98.0 98.0 98.0 100.0 100.0
imdb_reviews-em 96.0 96.0 98.0 98.0 98.0 98.0 98.0
sentiment140-em 68.0 68.0 68.0 66.0 66.0 68.0 70.0
READING Comp.
multirc-em 88.0 88.0 72.0 86.0 74.0 36.0 40.0
squad_v2-em 68.0 68.0 62.0 68.0 40.0 62.0 14.0
squad_v1-em 74.0 74.0 58.0 68.0 70.0 34.0 52.0
openbookqa-em 86.0 86.0 80.0 78.0 74.0 72.0 70.0
bool_q-em 84.0 84.0 68.0 82.0 76.0 66.0 80.0
drop-em 58.0 58.0 22.0 52.0 30.0 18.0 20.0
CLOSE-BOOK QA
natural_questions-em 30.0 30.0 28.0 24.0 18.0 12.0 14.0
arc_easy-em 94.0 94.0 90.0 92.0 84.0 88.0 86.0
arc_challenge-em 68.0 62.0 66.0 70.0 66.0 58.0 68.0
trivia_qa-em 68.0 70.0 56.0 58.0 64.0 54.0 58.0
COREFERENCE
definite_pronoun_resolution-em 88.0 88.0 90.0 88.0 74.0 72.0 62.0
wsc-em 62.0 64.0 60.0 64.0 64.0 58.0 48.0
READ. COMP. W/ COM
cosmos_qa-em 84.0 84.0 82.0 74.0 74.0 38.0 30.0
record-em 78.0 78.0 78.0 60.0 34.0 28.0 38.0
PARAPHRASE
paws_wiki-em 92.0 92.0 70.0 62.0 56.0 52.0 50.0
glue_qqp-em 86.0 86.0 88.0 82.0 84.0 76.0 80.0
glue_mrpc-em 84.0 84.0 70.0 80.0 70.0 62.0 62.0
stsb-em 44.0 44.0 44.0 44.0 28.0 20.0 22.0
NLI
cb-em 95.6 95.6 91.1 95.6 91.1 84.4 91.1
wnli-em 72.0 72.0 72.0 72.0 58.0 62.0 52.0
anli_r1-em 70.0 74.0 70.0 70.0 54.0 70.0 70.0
anli_r2-em 66.0 66.0 56.0 66.0 38.0 56.0 64.0
anli_r3-em 68.0 68.0 56.0 60.0 64.0 56.0 60.0
mnli_matched-em 86.0 88.0 90.0 88.0 88.0 88.0 88.0
mnli_mismatched-em 90.0 90.0 90.0 90.0 88.0 94.0 90.0
snli-em 90.0 90.0 88.0 90.0 82.0 88.0 88.0
qnli-em 94.0 94.0 94.0 84.0 66.0 30.0 58.0
rte-em 88.0 88.0 82.0 80.0 66.0 74.0 74.0
Table 18: Per-task results for the LoraRetriever weighted mixture routing method on LLaMA2-7B and LLaMA2-13B under Non-OOD and OOD evaluation.
Llama2-7B Llama2-13B
Domain-Metric Non-OOD OOD Non-OOD OOD
Struct to Text
web_nlg_en-rouge-1 60.9 48.5 68.6 52.2
web_nlg_en-rouge-2 36.7 26.5 47.2 28.5
web_nlg_en-rouge-l 54.0 43.8 61.6 46.7
dart-rouge-1 62.7 58.0 64.1 59.2
dart-rouge-2 36.7 33.4 39.9 36.1
dart-rouge-l 55.9 52.3 55.7 52.2
e2e_nlg-rouge-1 64.3 55.5 64.2 61.8
e2e_nlg-rouge-2 37.8 29.8 36.6 34.4
e2e_nlg-rouge-l 54.4 47.9 54.4 52.1
common_gen-rouge-1 50.1 29.8 46.1 28.2
common_gen-rouge-2 23.3 9.1 21.2 9.6
common_gen-rouge-l 46.5 26.2 42.6 24.4
Translation
para_crawl_enes-bleu 23.4 23.0 25.7 23.6
wmt16_translate_tren-bleu 3.2 2.6 2.5 2.8
wmt16_translate_deen-bleu 20.0 19.5 20.8 20.8
wmt16_translate_ruen-bleu 10.8 8.5 10.3 11.0
wmt16_translate_fien-bleu 6.5 7.1 6.8 8.3
wmt16_translate_roen-bleu 14.6 13.3 15.8 16.2
wmt14_enfr-bleu 17.3 18.2 15.7 18.8
wmt16_translate_csen-bleu 10.0 5.8 9.4 9.9
COMMONSENSE
story_cloze-em 80.0 88.0 96.0 94.0
piqa-em 42.0 36.0 52.0 48.0
copa-em 84.0 78.0 80.0 80.0
hellaswag-em 34.0 30.0 52.0 48.0
sentiment
sst2-em 98.0 94.0 98.0 100.0
yelp_polarity_reviews-em 98.0 98.0 100.0 100.0
imdb_reviews-em 98.0 100.0 98.0 98.0
sentiment140-em 68.0 72.0 72.0 72.0
READING Comp.
multirc-em 62.0 46.0 78.0 40.0
squad_v2-em 56.0 14.0 66.0 20.0
squad_v1-em 64.0 64.0 68.0 62.0
openbookqa-em 74.0 70.0 88.0 80.0
bool_q-em 80.0 78.0 84.0 86.0
drop-em 20.0 14.0 38.0 30.0
CLOSE-BOOK QA
natural_questions-em 12.0 10.0 28.0 18.0
arc_easy-em 74.0 80.0 94.0 90.0
arc_challenge-em 46.0 46.0 80.0 72.0
trivia_qa-em 56.0 50.0 62.0 56.0
COREFERENCE
definite_pronoun_resolution-em 60.0 64.0 92.0 68.0
wsc-em 56.0 50.0 70.0 56.0
READ. COMP. W/ COM
cosmos_qa-em 74.0 48.0 82.0 62.0
record-em 58.0 24.0 76.0 50.0
PARAPHRASE
paws_wiki-em 52.0 46.0 88.0 62.0
glue_qqp-em 86.0 60.0 82.0 82.0
glue_mrpc-em 62.0 64.0 72.0 64.0
stsb-em 50.0 12.0 44.0 24.0
NLI
cb-em 82.2 66.7 97.8 86.7
wnli-em 66.0 44.0 74.0 64.0
anli_r1-em 52.0 44.0 72.0 68.0
anli_r2-em 48.0 48.0 60.0 64.0
anli_r3-em 48.0 46.0 64.0 60.0
mnli_matched-em 84.0 80.0 86.0 84.0
mnli_mismatched-em 90.0 84.0 94.0 92.0
snli-em 86.0 84.0 92.0 88.0
qnli-em 80.0 74.0 80.0 72.0
rte-em 74.0 68.0 80.0 70.0